OpenStudy (anonymous):

Find three consecutive integers whose sum is 35

OpenStudy (niel):

11,12,13

OpenStudy (nikita2):

11 + 12 + 13 = 36

OpenStudy (nikita2):

May be you need solve next equation: x + x+1 + x+2 = 35, where x>=0 and integer.

OpenStudy (anonymous):

o.o but it's impossible. x can never be an integer

OpenStudy (nikita2):

Oh, yes)

myininaya (myininaya):

3x+3=35 3x=35-3 3x=32 x=32/3 gives not an integer is your question right? is it 36 or 35?

myininaya (myininaya):

nikita gj on finding the equation to use

myininaya (myininaya):

are you missing the word even or odd?

myininaya (myininaya):

let x be even then the next even integer is x+2 and next next even integer is x+4 so x+x+2+x+4=35 3x+6=35 3x=35-6 3x=29 x=29/3 not an integer so not even lets try odd for fun

OpenStudy (niel):

lol sorry im brainded today

myininaya (myininaya):

oh wait nvm it will the same way lol